# Nyström Kernel Stein Discrepancy

### v2

@@ 2 Notations
$C_{\mathbb{P},k}=\int_{\mathbb{R}^{d}}\phi_{k}(\mathbf{x})\otimes\phi_{k}(\mathbf{x})\mathrm{d}\mathbb{P}(\mathbf{x})\in\mathcal{H}_{k}\otimes\mathcal{H}_{k}$ (1)
exists if $\int_{\mathcal{X}}\left\|{\phi_{k}(\mathbf{x})}\right\|_{\mathcal{H}_{k}}^{2}\mathrm{d}\mathbb{P}(\mathbf{x})<\infty$; $C_{\mathbb{P},k}$ is a positive trace-class operator. We define $C_{\mathbb{P},k,\lambda}=C_{\mathbb{P},k}+\lambda I$, where $I$ denotes the identity operator and $\lambda>0$. The effective dimension of $\mathbb{P}\in\mathcal{M}_{1}^{+}\left(\mathbb{R}^{d}\right)$ is defined as $\mathcal{N}_{\mathbb{P},k}(\lambda):=\operatorname{tr}\left(C_{\mathbb{P},k}C_{\mathbb{P},k,\lambda}^{-1}\right)\leq\frac{\operatorname{tr}\left(C_{\mathbb{P},k}\right)}{\lambda}$. [^2] This inequality is implied by $\operatorname{tr}\left(C_{\mathbb{P},k}C_{\mathbb{P},k,\lambda}^{-1}\right)=\sum_{i\in I}\frac{\lambda_{i}}{\lambda_{i}+\lambda}\leq\frac{1}{\lambda}\sum_{i\in I}\lambda_{i}=\frac{\operatorname{tr}\left(C_{\mathbb{P},k}\right)}{\lambda}$, where $(\lambda_{i})_{i\in I}$ denote the eigenvalues of $C_{\mathbb{P},k}$. With $r\geq 1$ and a real-valued random variable $X:\left(\Omega,\mathcal{A},\mathbb{P}\right)\rightarrow\left(\mathbb{R},\mathcal{B}(\tau_{\mathbb{R}})\right)$, where $\mathcal{B}(\tau_{\mathbb{R}})$ denotes the Borel $\sigma$-field on $\mathbb{R}$, let $\left\|X\right\|_{L_{r}(\mathbb{P})}=\left[\int_{\Omega}|X(\omega)|^{r}\mathrm{d}\mathbb{P}(\omega)\right]^{\frac{1}{r}}$. For $r\in\{1,2\}$, let $\psi_{r}(u)=e^{u^{r}}-1$ and $\left\|{X}\right\|_{\psi_{r}}:=\inf\left\{C>0\mid\mathbb{E}_{X\sim\mathbb{P}}\psi_{r}\left(\frac{|X|}{C}\right)\leq 1\right\}$. A real-valued random variable $X\sim\mathbb{P}\in\mathcal{M}_{1}^{+}(\mathbb{R})$ is called sub-exponential if $\left\|{X}\right\|_{\psi_{1}}<\infty$ and sub-Gaussian if $\left\|{X}\right\|_{\psi_{2}}<\infty$. In the following, we specialize Definition 2 by Koltchinskii and Lounici [2017] stated for Banach spaces to (reproducing kernel) Hilbert spaces by using the Riesz representation theorem. A centered random variable $X\sim\mathbb{Q}\in\mathcal{M}_{1}^{+}\left(\mathcal{H}_{k}\right)$ taking values in an RKHS $\mathcal{H}_{k}$ is called sub-Gaussian iff. there exists a universal constant $C>0$ such that
$\left\|{\left\langle{X,u}\right\rangle_{\mathcal{H}_{k}}}\right\|_{\psi_{2}}\leq C\left\|{\left\langle{X,u}\right\rangle_{\mathcal{H}_{k}}}\right\|_{L_{2}(\mathbb{Q})}<\infty$ (2)
holds for all $u\in\mathcal{H}_{k}$.
@@ 3 Problem formulation
We now introduce our quantity of interest, the kernel Stein discrepancy. Let $\mathcal{H}_{k}^{d}:=\times_{i=1}^{d}\mathcal{H}_{k}$ be the product RKHS with inner product defined by $\left\langle{\mathbf{f},\mathbf{g}}\right\rangle_{\mathcal{H}_{k}^{d}}=\sum_{i=1}^{d}\left\langle{f_{i},g_{i}}\right\rangle_{\mathcal{H}_{k}}$ for $\mathbf{f}=\left(f_{i}\right)_{i=1}^{d},\mathbf{g}=\left(g_{i}\right)_{i=1}^{d}\in\mathcal{H}_{k}^{d}$. Let $\mathbb{P},\mathbb{Q}\in\mathcal{M}_{1}^{+}\left(\mathbb{R}^{d}\right)$ be fixed; we refer to $\mathbb{P}$ as the target distribution and to $\mathbb{Q}$ as the sampling distribution. Assume that $\mathbb{P}$ is absolutely continuous w.r.t. the Lebesgue measure $\lambda_{d}$ on $\mathbb{R}^{d}$; let the corresponding Radon-Nikodym derivative—in other words, the probability density function (pdf) of $\mathbb{P}$—be $p=\dfrac{\mathrm{d}\mathbb{P}}{\mathrm{d}\lambda_{d}}$. We assume that $p$ is continuously differentiable with support $\mathbb{R}^{d}$, $p(\mathbf{x})>0$ for all $\mathbf{x}\in\mathbb{R}^{d}$, and $\lim_{\left\|{\mathbf{x}}\right\|\to\infty}f(\mathbf{x})p(\mathbf{x})=0$ for all $f\in\mathcal{H}_{k}$. The last property holds for instance if $p$ is bounded and $\lim_{\left\|{\mathbf{x}}\right\|\to\infty}f(\mathbf{x})=0$ for all $f\in\mathcal{H}_{k}$. The Stein operator is defined as $\left(T_{p}\mathbf{f}\right)(\mathbf{x})=\langle\nabla_{\mathbf{x}}[\log p(\mathbf{x})],\mathbf{f}(\mathbf{x})\rangle+\langle\nabla_{\mathbf{x}}\mathbf{f}(\mathbf{x}),\mathbf{1}_{d}\rangle$ $\left(\mathbf{f}\in\mathcal{H}_{k}^{d},\,\mathbf{x}\in\mathbb{R}^{d}\right)$. With this notation at hand, for all $\mathbf{f}\in\mathcal{H}_{k}^{d}$ and $\mathbf{x}\in\mathbb{R}^{d}$ one has
$\left(T_{p}\mathbf{f}\right)(\mathbf{x})=\left\langle{\mathbf{f},\bm{\xi}_{p}(\mathbf{x})}\right\rangle_{\mathcal{H}_{k}^{d}},$ $\bm{\xi}_{p}(\mathbf{x})=\left[\nabla_{\mathbf{x}}\left(\log p(\mathbf{x})\right)\phi_{k}(\mathbf{x})+\nabla_{\mathbf{x}}\phi_{k}(\mathbf{x})\right]\in\mathcal{H}_{k}^{d},$ (3)
with the kernel
$h_{p}(\mathbf{x},\mathbf{y})$ $=\left\langle{\bm{\xi}_{p}(\mathbf{x}),\bm{\xi}_{p}(\mathbf{y})}\right\rangle_{\mathcal{H}_{k}^{d}}=\langle h_{p}(\cdot,\mathbf{x}),h_{p}(\cdot,\mathbf{y})\rangle_{\mathcal{H}_{h_{p}}},\quad\mathbf{x},\mathbf{y}\in\mathbb{R}^{d};$ (4)
@@
notice that $\bm{\xi}_{p}(\mathbf{x})$ and $h_{p}(\cdot,\mathbf{x})$ map to different feature spaces ($\mathcal{H}_{k}^{d}$ and $\mathcal{H}_{h_{p}}$, respectively) but yield the same kernel $h_{p}$. Throughout this work, we assume the measurability of $h_{p}$ and the separability of $\mathcal{H}_{h_{p}}$; this is satisfied, for example, if $k$ is continuously differentiable.
The kernel Stein discrepancy (KSD) [Chwialkowski et al., 2016, Liu et al., 2016] then is defined as
@@
$S_{p}(\mathbb{Q})$ $=\sup_{\mathbf{f}\in B\left(\mathcal{H}_{k}^{d}\right)}\left\langle{\mathbf{f},\mathbb{E}_{X\sim\mathbb{Q}}\bm{\xi}_{p}(X)}\right\rangle_{\mathcal{H}_{k}^{d}}=\left\|{\mathbb{E}_{X\sim\mathbb{Q}}\bm{\xi}_{p}(X)}\right\|_{\mathcal{H}_{k}^{d}}=\left\|{\mathbb{E}_{X\sim\mathbb{Q}}h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}},$ (5)
where the last equality follows from (4). By the construction of KSD,
$S_{p}(\mathbb{P})$ $=\left\|{\mathbb{E}_{X\sim\mathbb{P}}h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}}=0,$ (6)
that is, the expected value of the Stein feature map under the target distribution $\mathbb{P}$ is zero.
@@
Given a sample ^$\hat{\mathbb{Q}}_{n}=\{\mathbf{x}_{i}\}_{i=1}^{n}\sim\mathbb{Q}^{n}$, the popular V-statistic-based estimator [Chwialkowski et al., 2016, Section 2.2] is constructed by replacing $\mathbb{Q}$ with the empirical measure ^$\hat{\mathbb{Q}}_{n}$; it takes the form
@@
and can be computed in $\mathcal{O}\!\left(n^{2}\right)$ time. The corresponding U-statistic-based estimator [Liu et al., 2016, (14)] has a similar expression but omits the diagonal terms, that is, $S^{2}_{p,u}\left(\hat{\mathbb{Q}}_{n}\right)=\frac{1}{n(n-1)}\sum_{1\leq i\neq j\leq n}^{n}h_{p}(\mathbf{x}_{i},\mathbf{x}_{j})$; it also has a runtime costs of $\mathcal{O}\!\left(n^{2}\right)$. For large-scale applications, the quadratic runtime is a significant bottleneck; this is the shortcoming we tackle in the following.
@@ 4.1 The Nyström-KSD estimator
We consider a subsample ~$\tilde{\mathbb{Q}}_{m}=\left\{\left\{\tilde{\mathbf{x}}_{1},\ldots,\tilde{\mathbf{x}}_{m}\right\}\right\}$ of size $m$ (sampled with replacement), the so-called Nyström sample, of the original sample ^$\hat{\mathbb{Q}}_{n}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$; the tilde indicates a relabeling. The Nyström sample spans $\mathcal{H}_{p,m}=\mathrm{span}\left\{\bm{\xi}_{p}(\tilde{\mathbf{x}}_{i})\mid i\in[m]\right\}\subset\mathcal{H}_{k}^{d}$, with feature map $\bm{\xi}_{p}$ and associated kernel $h_{p}$ defined in (4).
One idea is to restrict the space over which one takes the supremum in (5), that is, one could consider
$S_{p}(\mathbb{Q})$ =(5)${{=}}\sup_{\mathbf{f}\in B\left(\mathcal{H}_{k}^{d}\right)}\left\langle{\mathbf{f},\mathbb{E}_{X\sim\mathbb{Q}}\bm{\xi}_{p}(X)}\right\rangle_{\mathcal{H}_{k}^{d}}\stackrel{{\scriptstyle(i)}}{{\approx}}\sup_{\mathbf{f}\in B\left(\mathcal{H}_{p,m}\right)}\left\langle{\mathbf{f},\mathbb{E}_{X\sim\mathbb{Q}}\bm{\xi}_{p}(X)}\right\rangle_{\mathcal{H}_{k}^{d}}$ (8)
$\stackrel{{\scriptstyle(ii)}}{{\approx}}\sup_{\mathbf{f}\in B\left(\mathcal{H}_{p,m}\right)}\left\langle{\mathbf{f},\mathbb{E}_{X\sim\hat{\mathbb{Q}}_{n}}\bm{\xi}_{p}(X)}\right\rangle_{\mathcal{H}_{k}^{d}}\stackrel{{\scriptstyle(iii)}}{{=}}\sqrt{\bm{\beta}_{p}^{\mkern-1.5mu\mathsf{T}}\mathbf{K}_{h_{p},m,m}^{-1}\bm{\beta}_{p}},$ (9)
with $\bm{\beta}_{p}=\frac{1}{n}\mathbf{K}_{h_{p},m,n}\bm{1}_{n}$. The details are as follows. In approximation $(i)$, we restrict the supremum to $\mathcal{H}_{p,m}\subset\mathcal{H}_{k}^{d}$, in step $(ii)$ the empirical measure ^$\hat{\mathbb{Q}}_{n}$ is used instead of $\mathbb{Q}$, and $(iii)$ is detailed in Appendix C. However, one can see that the intuitive estimator (9) has limited applicability, as the inverse of $\mathbf{K}_{h_{p},m,m}$ does not necessarily exist (for instance, by the non-strictly positive definiteness of $k$ or the sampling with replacement); hence, we take an alternative route which computationally leads to a similar expression and also allows statistical analysis.
The best approximation of $S_{p}(\mathbb{Q})$ in RKHS-norm-sense, when using $m$ Nyström samples, could be obtained by considering the orthogonal projection of $\mathbb{E}_{X\sim\mathbb{Q}}h_{p}\left(\cdot,{X}\right)$ onto $\mathcal{H}_{h_{p},m}:=\mathrm{span}\left\{h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)\mid i\in[m]\right\}\subset\mathcal{H}_{h_{p}}$. As $\mathbb{Q}$ is unknown in practice and only available via samples ^$\hat{\mathbb{Q}}_{n}\sim\mathbb{Q}^{n}$, we consider the orthogonal projection of $\mathbb{E}_{X\sim\hat{\mathbb{Q}}_{n}}h_{p}\left(\cdot,{X}\right)$ onto $\mathcal{H}_{h_{p},m}$ instead. In other words, we aim to find the weights $\bm{\alpha}=(\alpha_{i})_{i=1}^{m}\in\mathbb{R}^{m}$ that correspond to the minimum norm solution of the cost function
@@
$\tilde{S}_{p}^{2}\left(\hat{\mathbb{Q}}_{n}\right):=\left\|{\sum_{i=1}^{m}\alpha_{i}h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)}\right\|_{\mathcal{H}_{h_{p}},m}^{2}(=\left\|{P_{\mathcal{H}_{h_{p},m}}\mathbb{E}_{X\sim\hat{\mathbb{Q}}_{n}}h_{p}\left(\cdot,{X}\right))}\right\|_{\mathcal{H}_{h_{p},m}}^{2}).$ (11)
@@
where we define $\bm{\beta}_{p}=\frac{1}{n}\mathbf{K}_{h_{p},m,n}\bm{1}_{n}\in\mathbb{R}^{m}$, $\mathbf{K}_{h_{p},m,m}=\left[h_{p}\left(\tilde{\mathbf{x}}_{i},\tilde{\mathbf{x}}_{j}\right)\right]_{i,j=1}^{m}\in\mathbb{R}^{m\times m}$, and $\mathbf{K}_{h_{p},m,n}=\left[h_{p}\left(\tilde{\mathbf{x}}_{i},{\mathbf{x}}_{j}\right)\right]_{i,j=1}^{m,n}\in\mathbb{R}^{m\times n}$.
@@
    Comparison of (9) and (12). The two presented Nyström approaches result in comparable estimators — the inverse becomes a pseudo-inverse — but the projection perspective lends itself to a principled analysis that we detail in Section 4.3.

### v4

@@ 2 NOTATIONS
$C_{\mathbb{P},k}=\int_{\mathbb{R}^{d}}k\left(\cdot,\mathbf{x}\right)\otimes k\left(\cdot,\mathbf{x}\right)\mathrm{d}\mathbb{P}(\mathbf{x})\in\mathcal{H}_{k}\otimes\mathcal{H}_{k}$ (1)
exists if $\int_{\mathcal{X}}\left\|{k\left(\cdot,\mathbf{x}\right)}\right\|_{\mathcal{H}_{k}}^{2}\mathrm{d}\mathbb{P}(\mathbf{x})<\infty$; $C_{\mathbb{P},k}$ is a positive trace-class operator. We define $C_{\mathbb{P},k,\lambda}=C_{\mathbb{P},k}+\lambda I$, where $I$ denotes the identity operator and $\lambda>0$. The effective dimension of $\mathbb{P}\in\mathcal{M}_{1}^{+}\left(\mathbb{R}^{d}\right)$ is defined as $\mathcal{N}_{\mathbb{P},k}(\lambda):=\operatorname{tr}\left(C_{\mathbb{P},k}C_{\mathbb{P},k,\lambda}^{-1}\right)\leq\frac{\operatorname{tr}\left(C_{\mathbb{P},k}\right)}{\lambda}$.333 This inequality is implied by $\operatorname{tr}\left(C_{\mathbb{P},k}C_{\mathbb{P},k,\lambda}^{-1}\right)=\sum_{i\in I}\frac{\lambda_{i}}{\lambda_{i}+\lambda}\leq\frac{1}{\lambda}\sum_{i\in I}\lambda_{i}=\frac{\operatorname{tr}\left(C_{\mathbb{P},k}\right)}{\lambda}$, where $(\lambda_{i})_{i\in I}$ denote the eigenvalues of $C_{\mathbb{P},k}$. With $r\geq 1$ and a real-valued random variable $X:\left(\Omega,\mathcal{A},\mathbb{P}\right)\rightarrow\left(\mathbb{R},\mathcal{B}(\tau_{\mathbb{R}})\right)$, where $\mathcal{B}(\tau_{\mathbb{R}})$ denotes the Borel $\sigma$-field on $\mathbb{R}$, let $\left\|X\right\|_{L_{r}(\mathbb{P})}=\left[\int_{\Omega}|X(\omega)|^{r}\mathrm{d}\mathbb{P}(\omega)\right]^{\frac{1}{r}}$. For $r\in\{1,2\}$, let $\psi_{r}(u)=e^{u^{r}}-1$ and $\left\|{X}\right\|_{\psi_{r}}:=\inf\left\{C>0\mid\mathbb{E}_{X\sim\mathbb{P}}\psi_{r}\left(\frac{|X|}{C}\right)\leq 1\right\}$. A real-valued random variable $X\sim\mathbb{P}\in\mathcal{M}_{1}^{+}(\mathbb{R})$ is called sub-exponential if $\left\|{X}\right\|_{\psi_{1}}<\infty$ and sub-Gaussian if $\left\|{X}\right\|_{\psi_{2}}<\infty$. In the following, we specialize Definition 2 by Koltchinskii and Lounici [2017] stated for Banach spaces to (reproducing kernel) Hilbert spaces by using the Riesz representation theorem. A centered $\mathcal{H}_{k}$-valued random variable $X\sim\mathbb{Q}\in\mathcal{M}_{1}^{+}\left(\mathcal{H}_{k}\right)$ is called sub-Gaussian iff. there exists a universal constant $C>0$ such that for all $u\in\mathcal{H}_{k}$:
$\left\|{\left\langle{X,u}\right\rangle_{\mathcal{H}_{k}}}\right\|_{\psi_{2}}\leq C\left\|{\left\langle{X,u}\right\rangle_{\mathcal{H}_{k}}}\right\|_{L_{2}(\mathbb{Q})}<\infty.$ (2)
@@ 3 PROBLEM FORMULATION
We now introduce our quantity of interest, the kernel Stein discrepancy. Let $\mathcal{H}_{k}^{d}:=\times_{i=1}^{d}\mathcal{H}_{k}$ be the product RKHS with inner product defined by $\left\langle{\mathbf{f},\mathbf{g}}\right\rangle_{\mathcal{H}_{k}^{d}}=\sum_{i=1}^{d}\left\langle{f_{i},g_{i}}\right\rangle_{\mathcal{H}_{k}}$ for $\mathbf{f}=\left(f_{i}\right)_{i=1}^{d},\mathbf{g}=\left(g_{i}\right)_{i=1}^{d}\in\mathcal{H}_{k}^{d}$. Let $\mathbb{P},\mathbb{Q}\in\mathcal{M}_{1}^{+}\left(\mathbb{R}^{d}\right)$ be fixed; we refer to $\mathbb{P}$ as the target distribution and to $\mathbb{Q}$ as the sampling distribution. Assume that $\mathbb{P}$ is absolutely continuous w.r.t. the Lebesgue measure and let $p$ be the corresponding density (w.r.t. Lebesgue measure). We assume that $p$ is continuously differentiable with support $\mathbb{R}^{d}$, $p(\mathbf{x})>0$ for all $\mathbf{x}\in\mathbb{R}^{d}$, and $\lim_{\left\|{\mathbf{x}}\right\|\to\infty}f(\mathbf{x})p(\mathbf{x})=0$ for all $f\in\mathcal{H}_{k}$. The last property holds for instance if $p$ is bounded and $\lim_{\left\|{\mathbf{x}}\right\|\to\infty}f(\mathbf{x})=0$ for all $f\in\mathcal{H}_{k}$. Further, we assume that $k$ is continuously differentiable in both arguments. This condition will imply the measurability of $h_{p}$ and the separability of $\mathcal{H}_{h_{p}}$, both quantities defined below. The Stein operator [Gorham and Mackey, 2015, (4)] is defined as $\left(T_{p}\mathbf{f}\right)(\mathbf{x})=\langle\nabla_{\mathbf{x}}[\log p(\mathbf{x})],\mathbf{f}(\mathbf{x})\rangle+\sum_{i=1}^{d}\dfrac{\partial f_{i}(\mathbf{x})}{\partial x_{i}}$ $\left(\mathbf{f}\in\mathcal{H}_{k}^{d},\,\mathbf{x}\in\mathbb{R}^{d}\right)$. With this definition at hand,
$\left(T_{p}\mathbf{f}\right)(\mathbf{x})$ $=\left\langle{\mathbf{f},\bm{\xi}_{p}(\mathbf{x})}\right\rangle_{\mathcal{H}_{k}^{d}},$ (3)
$\bm{\xi}_{p}(\mathbf{x})$ $=\left[\nabla_{\mathbf{x}}\left(\log p(\mathbf{x})\right)k\left(\cdot,\mathbf{x}\right)+\nabla_{\mathbf{x}}k\left(\cdot,\mathbf{x}\right)\right]\in\mathcal{H}_{k}^{d}$ (4)
for all $\mathbf{f}\in\mathcal{H}_{k}^{d}$ and $\mathbf{x}\in\mathbb{R}^{d}$, with kernel (for $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$)
$h_{p}(\mathbf{x},\mathbf{y})$ $=\left\langle{\bm{\xi}_{p}(\mathbf{x}),\bm{\xi}_{p}(\mathbf{y})}\right\rangle_{\mathcal{H}_{k}^{d}}$ (5)
$=\langle h_{p}(\cdot,\mathbf{x}),h_{p}(\cdot,\mathbf{y})\rangle_{\mathcal{H}_{h_{p}}};$ (6)
@@
notice that $\bm{\xi}_{p}(\mathbf{x})$ and $h_{p}(\cdot,\mathbf{x})$ map to different feature spaces ($\mathcal{H}_{k}^{d}$ and $\mathcal{H}_{h_{p}}$, respectively) but yield the same kernel $h_{p}$, which, with (4), takes the explicit form
$h_{p}(\mathbf{x},\mathbf{y})$ $=\left\langle{\nabla_{\mathbf{x}}\log p(\mathbf{x}),\nabla_{\mathbf{y}}\log p(\mathbf{y})}\right\rangle_{\mathbb{R}^{d}}k(\mathbf{x},\mathbf{y})+\left\langle{\nabla_{\mathbf{y}}\log p(\mathbf{y}),\nabla_{\mathbf{x}}k(\mathbf{x},\mathbf{y})}\right\rangle_{\mathbb{R}^{d}}$ (7)
@@
$\quad+\left\langle{\nabla_{\mathbf{x}}\log p(\mathbf{x}),\nabla_{\mathbf{y}}k(\mathbf{x},\mathbf{y})}\right\rangle_{\mathbb{R}^{d}}+\sum_{i=1}^{d}\dfrac{\partial^{2}k(\mathbf{x},\mathbf{y})}{\partial x_{i}\partial y_{i}}.$
The kernel Stein discrepancy (KSD; Chwialkowski et al. 2016, Liu et al. 2016) then is defined as an integral probability metric [Zolotarev, 1983, Müller, 1997]
$S_{p}(\mathbb{Q})$ $=\sup_{\mathbf{f}\in B\left(\mathcal{H}_{k}^{d}\right)}\underbrace{\mathbb{E}_{X\sim\mathbb{P}}\left[T_{p}\mathbf{f}(X)\right]}_{\stackrel{{\scriptstyle(a)}}{{=}}0}-\;\mathbb{E}_{X\sim\mathbb{Q}}\left[T_{p}\mathbf{f}(X)\right]=\sup_{\mathbf{f}\in B\left(\mathcal{H}_{k}^{d}\right)}\left\langle{\mathbf{f},\mathbb{E}_{X\sim\mathbb{Q}}\bm{\xi}_{p}(X)}\right\rangle_{\mathcal{H}_{k}^{d}}$ (8)
$=\left\|{\mathbb{E}_{X\sim\mathbb{Q}}\bm{\xi}_{p}(X)}\right\|_{\mathcal{H}_{k}^{d}}\stackrel{{\scriptstyle(b)}}{{=}}\left\|{\mathbb{E}_{X\sim\mathbb{Q}}h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p}}},\;\;$ (9)
where (a) holds by the construction of KSD and (b) follows from (6).
@@
Given a sample ^$\hat{\mathbb{Q}}_{n}=\{\mathbf{x}_{i}\}_{i=1}^{n}\sim\mathbb{Q}^{n}$, the popular V-statistic-based estimator [Chwialkowski et al., 2016, Section 2.2] is obtained by replacing $\mathbb{Q}$ with the empirical measure ^$\hat{\mathbb{Q}}_{n}$; it takes the form
@@
and can be computed in $\mathcal{O}\!\left(n^{2}\right)$ time. The corresponding U-statistic-based estimator [Liu et al., 2016, (14)] has a similar expression but omits the diagonal terms, that is,
$S^{2}_{p,u}\left(\hat{\mathbb{Q}}_{n}\right)=\frac{1}{n(n-1)}\sum_{1\leq i\neq j\leq n}^{n}h_{p}(\mathbf{x}_{i},\mathbf{x}_{j});$ (11)
it also has a runtime cost of $\mathcal{O}\!\left(n^{2}\right)$. For large-scale applications, the quadratic runtime is a significant bottleneck; this is the shortcoming we tackle in the following.
@@ 4.1 The Nyström-KSD Estimator
We consider a subsample ~$\tilde{\mathbb{Q}}_{m}=\left\{\left\{\tilde{\mathbf{x}}_{1},\ldots,\tilde{\mathbf{x}}_{m}\right\}\right\}$ of size $m$ (sampled with replacement), the so-called Nyström sample, of the original sample ^$\hat{\mathbb{Q}}_{n}=\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\right\}$; the tilde indicates a relabeling. The best approximation of $S_{p}(\mathbb{Q})$ in RKHS-norm-sense, when using $m$ Nyström samples, can be obtained by considering the orthogonal projection of $\mathbb{E}_{X\sim\mathbb{Q}}h_{p}\left(\cdot,{X}\right)$ onto $\mathcal{H}_{h_{p},m}:=\mathrm{span}\left\{h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)\mid i\in[m]\right\}\subset\mathcal{H}_{h_{p}}$, with feature map $h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)$ and associated kernel $h_{p}$ defined in (6). As $\mathbb{Q}$ is unknown in practice and only available via samples ^$\hat{\mathbb{Q}}_{n}\sim\mathbb{Q}^{n}$, we consider the orthogonal projection of $\mathbb{E}_{X\sim\hat{\mathbb{Q}}_{n}}h_{p}\left(\cdot,{X}\right)$ onto $\mathcal{H}_{h_{p},m}$ instead. In other words, we aim to find the weights $\bm{\alpha}=(\alpha_{i})_{i=1}^{m}\in\mathbb{R}^{m}$ that correspond to the minimum norm solution of the cost function
@@
$\tilde{S}_{p}^{2}\left(\hat{\mathbb{Q}}_{n}\right)$ $:=\left\|{\sum_{i=1}^{m}\alpha_{i}h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)}\right\|_{\mathcal{H}_{h_{p}},m}^{2}=\left\|{P_{\mathcal{H}_{h_{p},m}}\mathbb{E}_{X\sim\hat{\mathbb{Q}}_{n}}h_{p}\left(\cdot,{X}\right)}\right\|_{\mathcal{H}_{h_{p},m}}^{2}.$ (13)
@@
where $\bm{\beta}_{p}=\frac{1}{n}\mathbf{K}_{h_{p},m,n}\bm{1}_{n}\in\mathbb{R}^{m}$, Gram matrix $\mathbf{K}_{h_{p},m,m}=\left[h_{p}\left(\tilde{\mathbf{x}}_{i},\tilde{\mathbf{x}}_{j}\right)\right]_{i,j=1}^{m}\in\mathbb{R}^{m\times m}$, and $\mathbf{K}_{h_{p},m,n}=\left[h_{p}\left(\tilde{\mathbf{x}}_{i},{\mathbf{x}}_{j}\right)\right]_{i,j=1}^{m,n}\in\mathbb{R}^{m\times n}$.
@@
    Comparison to Chatalic et al. [2022]. We note that the estimator (14) corresponds precisely to Chatalic et al. [2022, (5)]. We consider the analysis of this known estimator in the case of unbounded feature maps—which arise in the KSD setting—as one of our core contributions, which we detail in Section 4.3.